in a school 3/5 of the students are girls. if the number of boys in the school is 200 , find the total number of students in school
Answer by rothauserc(4718)

if 3/5 are girls then 2/5 are boys, then let x be the total number of students
:
(2/5) * x = 200
:
2x = 1000
:
x = 500
:
The total number of students in school is 500
:
Note that (3/5) * 500 = 300 and 300 + 200 = 500
: